CTIO
Published on CTIO (http://www.ctio.noao.edu/noao)

CTIO Home > Optical Engineering > Maintenance Procedures > Thermal issues

Thermal issues

  •  Where to find useful logs? Temp4m logs are stored in /ut02/seeing/temper/data/. Weather logs are stored in /ut02/seeing/weather_data/ (data prior to 1998 is still lost...). Site seeing logs are stored in /ut02/seeing/seeing_data/. The focus log at PF is /ut02/seeing/focuslog. The old 4m tel side port seeing camera log is /ut02/seeing/jacoby.log.
  • Memo (18Oct99 [1]) describing the revised operation instructions at the 4m for image quality program (includes seeing measurements, thermal environment,...)
  • Is the mirror going to be wet? [2] (check the risk of condensation)
  • Primary mirror cooling (active only during day-time, at night time we simply reverse the system and extract air to prevent hot air bubles to form above the mirror). The 15 tons of Cervit have a large thermal inertia, but we can typically cool its surface by 0.75deg/hour by blowing 6deg cooler air around it. The control algorithm is rather simple as it is based on the fact that usually the dome temperature will reflect well the variations of the outside temperature, thus by keeping the mirror 2deg colder than the low dome during the day, we start the next night in optimum condition. The typical night time pattern is for the outside temperature to drop slowly and cross the mirror temperature at about 2/3 of the night to finish some 0.5deg colder at morning twilight. This graph [3]shows statistics of the difference between mirror and low dome temperatures.
  • Temperature sensors [4] (sensor list and some debugging hints for Temp4m)
  • Last improvements in the 4m dome for optimizing thermal environment:

    • The old console room walls were removed to open the direct pass of air flow blowing from the lateral doors into the dome (whenever the dome azimut is such that they are behind the console). These walls were 10m long. (March 2000)
    • The old mirror cover made of 24 petals opening up and forming a cylinder about 1.6m high (thus putting the mirror surface at the bottom of a 3.5m 'hole') was replaced by a 2-petal design (see picture 1 [5] and picture 2 [6]) opening up on the west and east side (just in front of the top ends of the horseshoe). This leaves free path to the air flow along the north-south direction (at zenith) and provides better (closer) ventilation to the primary mirror. See picture. (August 2000)
    • The dome is being covered with aluminium sticky foil [7] (September-November 2000).

The new skin of the 4.0-m dome

Up to a few years ago, there were traditionally 2 kind of paints used to cover telescope domes: 'Lomit' and titanium oxide.

Lomit is the grey-looking (silver- or aluminium-like) paint. It typically overheats by up to 20degC during the day in the sunshine and overcools by about 2degC at night after reaching equilibrium with ambient air (delay due to thermal inertia of structure).Its emissivity is 0.21.

Titanium oxide is the white paint which had, for example, always been in use at CTIO. It typically maintain thermal equilibrium with ambient air during the day (no overheat) but overcools by about 5degC during the night. Thus, although titanium oxide is far better than Lomit during the day, it is slightly worse at night, and the air cooling around the dome tends to fall/roll down off the walls and eventually enter the dome through the shutter or the lateral doors, eventually causing dome seeing. It is frequent in summer time on Cerro Tololo to see the domes wet while the RH is only 75%: this is because the outside skin of the domes overcools several degrees below ambient and can drop below the dew point.

In the early nineties, our colleagues at Las Campanas Observatory have used an aluminium sticky foil to cover their domes and got very positive results (more recently they used it for the Magellan twin telescope domes). Our measurements have shown that, with such a coating, the dome typically overheats by 5degC during the day and by about 2degC during the night. It has therefore better characteristics than the 2 older techniques.

Summary:

  • white titanium oxide paint: day=ambient ; night=ambient-5deg
  • grey Lomit paint: day=ambient+20deg ; night=ambient-2deg
  • aluminium foil: day=ambient+5deg ; night=ambient+2deg

The difference with the Lomit paint (both are greyish and reflectivish) is mainly that the foil is made out of an insulating layer of adhesive about 0.12mm thick (5/1000") coated with aluminium. Consequently there is no direct thermal contact between the ambient air and the metallic structure of the dome (thus avoiding the diurnal overheating). The product is called Compaq #804 Aluminium foil tape, it comes in rolls of 30"x60yds and it is sold by Bron tapes Inc. Adhesion is increased by application of a primer on the substrate, and is good enough that no maintenance is required over periods as long as 7 years. The foil can also be washed without deterioration.

Our hope is therefore to improve the thermal performance of the 4m dome and decrease dome seeing. New temperature sensors with RF transmitters and receivers will be mounted on the dome (which is a rotating part!) to attempt quantifying these modifications.

 

 

M.B., 14th November 2000

Thermal issues (hot stuff...)

Last updated Sept 1st, 2000

TEMPERATURE SENSORS LOCATION AND LIST

See drawing CH2917.C001

On February 9th, we have replaced the old temp4m PC with the new Labview-based system. The data is accessed by typing 'temp4m' from any machine. That will show you the important fields logged in /ut02/seeing/temper/newtemp.log which are described below:

 

1 2 3=A 4=B 5=C 6=D 7=E 8=F 9=G 10=H
UT time UT date Lo tru NW Lo tru NE Lo tru E Lo tru W  Up tru N Up tru E Up tru S Up tru W
11=I 12=J 13=K 14=L       15=M 16=N 17=O 18=P 19=Q 20=R
unused Mirror glycol Mc air out Oil pump Low dome Oil bef pad Oil aft dome High dome PF cage Cass cage out
21=S 22=T 23=U 24=V 25=W 26=X 27=Y 28=Z 29=a 30=b
Mirror W Chimney Mirror N Mirror E Mirror S outside Dome floor Oil glycol Dome RH Oil volt read
31=c 32=d 33=e 34=f 35=g 36=h 37 38 39 40
MC volt read Oil pump status Cass cafe in air M1 cell air above M1 guider air dew point oil volt calc mc volt calc  

 

Notes:

  •  the letters are the addresses of the DGHs
  • Mc stands for "mirror cooler"
  • 35 and 36 added on 31Aug00
  • Since 04/nov/99, both temperature logs (the old one 'temp.log' and the new one) read UT time (instead of local time before).

 

SOME QUICK CHECKS THAT THE THERMAL SYSTEM IS WORKING!!

Typing "temp4m" on a Sun machine gives a list of temperature:

See drawing CH2890-A001 for mirror cooler details.

 

MIRROR COOLER:

  • In day time: "mc glycol" should be negative typically between -5° and 0°(if the cooler works)
  • At night time: "mc glycol" gently comes back to ambient temperature
  • In day time: "Valve voltage" should read between 0 and 5 volts.
  • At night time (cooling off): a voltage is still calculated by the PC and logged but there is a pressure gauge in the air duct (just after the cooler) which, after the sucker mode is being selected at the start of the night, detects 0 pressure in the lower air duct and will override the PC voltage and send 0 to the Mamac so that the valve is closed. The glycol stops flowing and warms up slowly to ambient during the night (#5 in temp.log).

 

OIL:

  • Day and night: "Glycol" should be negative typically between -5° and 0°. If this is not the case, think about an eventual problem in the glycol room (tank + pumps) located in the building next to Las Tacas. When glycol doesn't flow to cool the oil, "Before Pad" and "Glycol" will say typically +30° (oil warms up to ambient temperature and a bit higher at night when the telescope moves). During day time the oil valve is closed and the "oil glycol" sensor will be a good indicator of the glycol temperature sent by the plant and getting to the mirror cooler system (which comes second in the system after the oil loop).
  • At night time: "Valve voltage" should be >0 and "Pump On/Off" should say 1 (if telescope is active)
  • In day time: "Valve voltage" should be =0 and "Pump On/Off" should 0

 

DOME and TEL:

  • Day and night, "Low dome", "High dome", "Serr truss", "Primary North", South" and "West" should always be within 1-2° from each other (and closer at night time), the mirror temperatures being hopefully slightly lower than the dome ones.
  • Cass cage is always a few degrees hotter than the previous set of sensors.

 

GLYCOL SYSTEM:

  • There are 2 separate circuits starting from the Glycol room: one for the building air conditioning and the other for the 4m telescope of : oil + mirror air cooler + M/C floors.

 

Notes about the old 'temp.log':

a- On 04/Nov/99, we found out errors in field labels in temp.log:

* #3: previously saying south, was actually east
* #4: previously saying west, was actually south
* #6: previously saying north, was actually west

b- On 04/nov/99, we modified the labels on the old temp4m PC display and swap connectors at the multiplexer input in order to have the PC algorithm use the correct sensors and log a trustful temp.log as indicated in the table above.

c- the old grey relay box in the Cass cage has the following inputs for the mirror sensors: #1 with cable labeled north, #2 with east, #3 with south and #4 with west. DO NOT change that!

d- the values displayed by the multiplexer in Tolnet show small variations (+/- 1°) from the reality (due to sensor calibration, effects of the cables,...) and serves only for debugging. The old PC algorithm uses a correction table for each sensor (file "tcoef").


Source URL (retrieved on 02/09/2013 - 09:15): http://www.ctio.noao.edu/noao/content/Thermal-issues

Links:
[1] http://www.ctio.noao.edu/noao/content/thermal-control-procedures
[2] http://www.ctio.noao.edu/noao/content/relative-humidity
[3] http://www.ctio.noao.edu/noao/sites/default/files/telescopes/mir_lodome.gif
[4] http://www.ctio.noao.edu/noao/content/Thermal-issues-hot-stuff
[5] http://www.ctio.noao.edu/noao/sites/default/files/telescopes/2petal1.jpg
[6] http://www.ctio.noao.edu/noao/sites/default/files/telescopes/2petal2.jpg
[7] http://www.ctio.noao.edu/noao/content/new-skin-40-m-dome